OpenStudy (anonymous):

Which is the equation to the line passing through points (2, -7) and (5, -1)?

OpenStudy (anonymous):

y = 2x - 8 y = - 2x + 9 y = 2x - 11 y = - 2x + 6

OpenStudy (anonymous):

those are the answers.

hartnn (hartnn):

do u know how to find slope when 2 points are given ?

OpenStudy (anonymous):

No

hartnn (hartnn):

The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

hartnn (hartnn):

could u find m=... ?

OpenStudy (anonymous):

No, im having a difficult time, i think its due to the early-ness. haha

hartnn (hartnn):

\(\huge m=\frac{y_1-y_2}{x_1-x_2}=\frac{-7-(-1)}{2-5}=...?\)

OpenStudy (anonymous):

-6/-3

hartnn (hartnn):

m= -6/-3 = ..... ?

OpenStudy (kainui):

It might be easier if you have a graphing calculator. Make a line that looks like: y=(2/3)x and then change the slope around a bit to figure it out and see why the slope is rise/run by going to the "trace" option on your calculator and trying things out, like when x=3.

OpenStudy (anonymous):

2

hartnn (hartnn):

yes, m=2

OpenStudy (compassionate):

Now use the first points given and the slope. Plug it into the equation: y - y1 = m(x - x1)

hartnn (hartnn):

now take any point, say (5,-1) and put it in y - y1 = m(x - x1)

hartnn (hartnn):

x1=5,m=2,y1=-1 isolate y

OpenStudy (anonymous):

-_-

OpenStudy (compassionate):

Tina, let me make it easier for you. y - y1 = m(x - x1) y - (-1) = 2(x - 5) <-- Solve for y

OpenStudy (anonymous):

Good lord this is terrible. I feel so confused.

OpenStudy (compassionate):

Okay. What are you confused with?

OpenStudy (anonymous):

Nothing. I think i'm going to skip this! I have to have it done by later tonight. I think im going to rest now haha thanks though:)

OpenStudy (compassionate):

Bye.